40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5061     Entry time: Fri Jul 29 16:04:51 2011
Author: steve 
Type: Update 
Category: SUS 
Subject: ETMX, PRM & SRM OSEM gains set 

Jamei fixed the computer. Now I had a healthier ETMX with SIDE signal that allowed me to set gains to get Q of 5

The Y arm restore locked the arm at TRY -out 2.4V,    the X arm still did not lock or align to lock. It has some fringes.

ETMX  POS gain   22  -> 17,              PIT  3  ->  1,                    YAW  3,                     SIDE     -170    ->    -110

PRM                        50  ->  30,                     1,                                       2,                                       50

SRM                        55  ->  20,                    2.4  ->  1.5,                      4.8  ->  2,                          140   ->    100

Now all SUS OSEM gains are set

ELOG V3.1.3-